OpenStudy (anonymous):

(Pre-cal: law of cosines) Find the area of the triangle. A=47 degrees b=32ft c=19 ft @amistre64

OpenStudy (amistre64):

law of cosine is a general version of the pythag thrm

OpenStudy (amistre64):

in this case: a^2 = b^2 + c^2 - 2bc cosA

OpenStudy (amistre64):

when the angle given is 90 degrees this simplifies to: a^2 = b^2 + c^2 such that a is the hypotenuse

OpenStudy (anonymous):

Oh, so I need to find a before I can do s=a+b+c/2 and plug in for are.

OpenStudy (anonymous):

area*

OpenStudy (amistre64):



OpenStudy (anonymous):

Then plug in s and the values of a, b, and c in the area formula. A=sqrts(s-a)(s-b)(s-c)

OpenStudy (amistre64):

if you want to do herons formula then yes, you would need to know all the side lengths

OpenStudy (anonymous):

Okay, well that's the only way I learned so I'm going to do it that way. But thank you! I needed a refresher :)

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks. :)